The replacement mirror for APASS-north has arrived and will be installed shortly. Of course, the southwestern monsoon is still producing clouds, so it may be a couple of weeks before we start getting new data.
APASS-south is broken again. There is some continuing problem with the filter wheel. The technician will look at it soon; hopefully it will be something easy to fix. We did get a few nights before things quit, and the replacement camera seems to be working quite well. Kevin Ivarsen has gotten the webcam working again, so we can watch the telescope when necessary. We have extended the monitoring queue to include T Pyx, PR Lup, N Sco 2011#2, and ASAS 105134 (for Sebastian Otero). Those stars are primarily monitored during bright time; we can add a few more if researchers have some interesting targets in the 9<V<14 range.